OpenStudy (anonymous):

Sheryl's salary is $1200 per month plus a 10% commission on her sales. What amount of sales would she need to earn at least $2000 per month? a. Her sales must be over $8000. b. Her sales must be $8000. c Her sales must be at least $8000. d. her sales must be less than $8000.

OpenStudy (anonymous):

B...She must sell $8000 ;)

OpenStudy (anonymous):

let her sale x \[1200+0.1x \ge 2000\] so 0.1x\[0.1x \ge 800, x \ge 8000\]